OpenStudy (anonymous):

implicit differentiate -y/(x-y)=x^(9)+3

OpenStudy (bahrom7893):

ill get to this in 5 mins..

OpenStudy (anonymous):

anyone know how to solve it when it gets to y'(x-2y)-y(1-2y')=9x^(8)(x-2y)^(2) ?

OpenStudy (bahrom7893):

write y' as dy/dx

OpenStudy (anonymous):

i know but it's much easier to acknowledge as y'

OpenStudy (bahrom7893):

it actually isnt.. u need to solve the ODE right?

OpenStudy (anonymous):

ODE?

OpenStudy (bahrom7893):

ordinary differential equation?

OpenStudy (anonymous):

are you in calc 1 or differential eq

OpenStudy (anonymous):

Calculus 1

OpenStudy (anonymous):

then it is probably not ode

OpenStudy (anonymous):

yes; -y = x^10 + 3x - yx^9 -3y -y' = 10x^9 + 3 - 9yx^8 - y'x^9 -3y' y'(x^9 + 2) = 10x^9 + 3 - 9yx^8 y' = (10x^9 + 3 + 9yx^8)/(x^9+2)

OpenStudy (bahrom7893):

we did ode in calc 1